A large population consist of equal numbers of the digits 1 and 3.

Find the mean and variance of this population


1
Expert's answer
2021-09-21T02:22:43-0400

We have here: 

P(X = 1) = P(X = 3) = 0.5

Therefore, Mean = E(X) = (1 + 3) / 2 = 2 is the required mean value here. 

E(X2) = (12 + 32) / 2 = 5

Var(X) = E(X2 ) - Mean2 = 5 - 22 = 1 is the required variance here.